### v1

# Towards a Quantum Information Theory of Hadronization: Dihadron Fragmentation and Neutral Polarization in Heavy Baryons

Abstract
We pioneer the application of quantum information theory to experimentally distinguish between classes of hadronization models. We adapt the CHSH inequality to the fragmentation of a single parton to hadron pairs, a violation of which would rule out classical dynamics of hadronization altogether. Furthermore, we apply and extend the theory of quantum contextuality and local quantum systems to the neutral polarization of a single spin-1 hadronic system, namely the light constituents of excited Sigma baryons $\Sigma^{*}_{c,b}$ formed in the fragmentation of heavy quarks.
preprint: MIT-CTP 5818preprint: Nikhef 2024-020preprint: DESY-25-034
I Introduction
Every single time that quarks and gluons are produced in particle collisions, they combine back into color-neutral bound states like the proton or the pion, collectively called hadrons. Calculating this nonperturbative, strongly coupled phenomenon called hadronization from first principles using the Lagrangian of Quantum Chromodynamics (QCD) remains an open problem, restricting us to phenomenological models. Perhaps surprisingly, and despite the underlying dynamics of quark and gluons clearly being quantum mechanical (QM) in general, classical stochastic models of hadronization Field and Feynman (1978); Field (1989); Andersson et al. (1983); Webber (1984) as e.g. implemented in multi-purpose Monte-Carlo (MC) generators Bierlich et al. (2022); Bewick et al. (2024); Bothmann et al. (2024) are vastly successful in this regard, notably for open, semi-inclusive hadronization observables. A macroscopic, statistical-mechanics approach especially succeeds in the large-multiplicity environment of heavy-ion collisions Hagedorn (1965); Becattini and Heinz (1997); Andronic et al. (2003); Chojnacki et al. (2012). In this letter we aim to identify experimental measurement outcomes that indicate a departure from classical expectations and – if observed – can only be described by a full QM treatment of hadronization, see e.g. Refs. Metz (2002); Kerbizi and Lönnblad (2022, 2023); Masouminia and Richardson (2024). This, to the reader, will be a familiar task: Asking whether local reality itself is quantum in the form of Bell’s inequality Bell (1964) lies at the root of today’s rich field of quantum information (QI) theory. We will therefore adapt and extend QI concepts to hadronization physics to ask, broadly speaking: Is hadronization quantum?
Our proposal is part of a larger fruitful effort to bring QI theory to bear on questions in particle physics Tornqvist (1981); Ablikim et al. (2019); Severi et al. (2022); Aoude et al. (2022); Afik and de Nova (2023); Aad et al. (2024); Barr et al. (2024); Wu et al. (2024a); Morales (2024); Chen et al. (2024); Kowalska and Sessolo (2024); Fabbrichesi and Marzola (2024); Bernal et al. (2024); Hayrapetyan et al. (2024); Afik et al. (2024); Guo et al. (2025); White and White (2024); Wu et al. (2024b); Fabbrichesi et al. (2024); Demina and Landi (2025); Ruzi et al. (2024); Morales and Szynkman (2024); Grossi et al. (2024); Cheng et al. (2024); Fabbrichesi et al. (2025a); Cheng and Yan (2025); Fabbrichesi et al. (2025b, c), but distinct in that most of the existing work, if not attempting to explain the Standard Model (SM) Mulders (2018); Thaler and Trifinopoulos (2024), aims to probe quantum foundations either at the highest energy scales or in exclusive hadronic processes, essentially asserting the quantum nature of the perturbative SM Lagrangian or hadronic form factors, respectively. Hadronization then either serves as an analyzer for the perturbative density matrix Aidala and Rogers (2021); Afik et al. (2024); Cheng and Yan (2025) or, possibly, as a source of decoherence Gong et al. (2022); Datta et al. (2024) acting on the entangled state created during the perturbative scattering. By contrast, our proposal directly concerns the nature of hadronization itself, and we appeal to principles of effective field theory and factorization theorems to isolate the nonperturbative behavior under study from the (in our view, most likely quantum, and calculable) perturbative scattering and the self-analyzing hadronic decays. Our work is also distinct from proposals to simulate hadronization on quantum circuits Florio et al. (2023); Bauer et al. (2023); Florio et al. (2024); Li et al. (2024); Lee et al. (2024); Bauer (2025).
II Classical hadronization models as “hidden” variables
In the picture we are developing, Bell’s “hidden variables” $\lambda$ take on concrete meaning as a point in the phase space of a stochastic hadronization model with an intermediate classical state $p(\lambda)$, i.e., a probability density. Of course, in this case the variables are not really hidden: In an analytic model, one may calculate their statistical distribution; if the model is implicitly defined by a stochastic algorithm, e.g. a Markov process acting on an ensemble of strings or preconfined clusters Amati and Veneziano (1979); Andersson et al. (1983); Webber (1984), one can print out and bin the intermediate state to record $p(\lambda)$. Just as in Bell’s concise proof Bell and Aspect (2004) of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al. (1969), suitable correlation observables $\mathcal{C}$ will have an expectation value $\langle\mathcal{C}\rangle_{p(\lambda)}\leq\mathcal{C}_{\mathrm{max}}$ that is bounded when computed in any classical hadronization model, following basic properties of the integral measure in any number of classical phase-space dimensions, while QM expectation values can exceed the bound. Note that $p(\lambda)$ and the conditional measurement probabilities of classical hadronization models must obey the spacetime symmetries of QCD (Lorentz covariance and parity), a constraint that is typically absent in tests of quantum foundations, but which we will make productive use of in this letter.
To fix ideas, consider the fragmentation $i\to h_{1}h_{2}X$ of an unpolarized parton $i=q,g$ into identified hadrons $h_{1,2}$. If $h_{1,2}$ are spin-$1/2$ baryons, their most general spin-density matrix compatible with the symmetries of QCD reads
$\rho=\frac{\mathbb{1}}{4}+\tilde{D}_{LL}\,\hat{S}_{1,L}\otimes\hat{S}_{2,L}+\tilde{D}_{TT}\,\hat{\mathbf{S}}_{1,T}\otimes\hat{\mathbf{S}}_{2,T}\,,$ (1)
where $\hbar=1$ and $\hat{S}_{j,L}$ ($\hat{\mathbf{S}}_{j,T}$) is the longitudinal (transverse) component of the spin operator of hadron $h_{j}$ with respect to the fragmentation axis $\mathbf{z}$ pointing back to the hard collision in the hadron’s rest frame. (In the last term, a contraction between vector indices is understood.) Eq. (1) is the polarization state of the system at long distances of order the hadron lifetimes, $t_{\mathrm{decay}}\sim 1/\Gamma_{h}$, where a QM description of the system is appropriate for the exclusive hadronic decays.111 Decay matrix elements and differential distributions are collected in the supplemental material. By contrast, the nonperturbative hadronization physics are encoded in normalized dihadron fragmentation functions (DiFFs) Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025) $\tilde{D}_{LL}$ Huang et al. (2025) and $\tilde{D}_{TT}$, which are determined from experiments or computed in a model. In the latter case, we interpret the DiFFs as matching coefficients that are fixed by computing spin expectation values $T_{k\ell}/4\equiv\langle S_{1,k}\otimes S_{2,\ell}\rangle$ in a (potentially classical) hadronization model at the scale $t_{\mathrm{confine}}\sim 1/\Lambda_{\mathrm{QCD}}\ll t_{\mathrm{decay}}$ where the hadrons form and decouple. We also have $t_{\mathrm{confine}}\gg t_{\mathrm{hard}}$, where $t_{\mathrm{hard}}$ is the very short time scale of the perturbative hard scattering. This separation of all scales allows us to specifically probe the quantum nature of hadronization and, while abstract, mirrors the workflow of MC generators.
The DiFFs completely characterize the system, see Fig. 1: Positivity requires $\tilde{D}_{LL}\geq-1$ and $\tilde{D}_{LL}+2\lvert\tilde{D}_{TT}\rvert\leq 1$. By the positive-partial-transpose (PPT) criterion Horodecki et al. (1996), the hadrons are entangled if and only if $\lvert\tilde{D}_{TT}\rvert>(\tilde{D}_{LL}-1)/2$. Most interestingly, we can evaluate the optimal Horodecki et al. (1995) CHSH inequality Clauser et al. (1969) for Eq. (1) in terms of the DiFFs,
$\tilde{D}_{TT}^{2}+\max\{\tilde{D}_{LL}^{2},\tilde{D}_{TT}^{2}\}\stackrel{{\scriptstyle p(\lambda)}}{{\leq}}1\,.$ (2)
This is the key result for this example process, illustrating our proposal: If Eq. (2) is found to be violated in data, the observation cannot be explained by any classical hadronization model. The proof exactly follows Ref. Bell and Aspect (2004), with one crucial difference: In our case, the “locality” assumption that measurements on $h_{1}$ are independent of the reconstruction of $h_{2}$ is guaranteed by color confinement, rather than by hypothesizing, like Bell, a local hidden-variable theory (LHVT) of Nature. The CHSH test we propose here therefore does not test quantum nonlocality in the usual way, but instead uses the identical mathematics to test whether the blackbox Hamiltonian of hadronization admits a classical description.222 Another assumption is that Gedanken measurements $\langle(\bm{e}_{\alpha}\cdot\bm{S}_{1})\otimes(\bm{S}_{2}\cdot\bm{e}_{\beta})\rangle$ at Bell angles $\alpha,\beta$ are indeed predicted to evaluate to $\bm{e}_{\alpha}\cdot(\mathsf{T}/4)\cdot\bm{e}_{\beta}$, where $\mathsf{T}=(T_{k\ell})$. In our case, this property is inherited from the interpolating partonic spin operators defined in the underlying QCD theory. Therefore, the counterexamples of Refs. Kasday et al. (1970); Abel et al. (1992) (nota bene constructed to argue against the viability of testing LHVTs of Nature at colliders) do not apply.
Figure 1: QI characterization of parton fragmentation to two nearby polarized hadrons. An observation of dihadron fragmentation functions taking values in the blue region is inconsistent with classical models of hadronization. For $\tilde{D}_{LL}=-1$ and $\tilde{D}_{TT}=-1$ ($+1$), the hadrons are produced in a pure singlet (triplet) Bell state.
III Quantum contextuality and local spin-1 systems
A fundamental aspect of QM systems is their (statistical) contextuality Budroni et al. (2022), the idea that a measurement’s outcome may depend on which other properties are being measured. To adapt the notion of statistical contextuality to classical hadronization models, consider a very simple hadronizing system: the production of excited heavy spin-$3/2$ baryons $\Sigma_{Q}^{*}$ during the fragmentation $Q\to\Sigma_{Q}^{*}X$ of beauty or charm quarks, $Q=b,c$. The light hadron constituents that bind to the static heavy quark carry total angular momentum $j=1$, forming a qutrit. By parity and rotational invariance about the fragmentation axis $\mathbf{z}$, their quantum polarization state $\rho$ at $t_{\mathrm{decay}}\sim 1/\Gamma_{\Sigma_{Q}^{*}}\gg t_{\mathrm{confine}}$ is governed by a single nonperturbative Falk-Peskin parameter $w_{1}$ Falk and Peskin (1994),
$\rho=\frac{w_{1}}{2}\bigl{(}\lvert 1_{z}\rangle\langle 1_{z}\rvert+\lvert-1_{z}\rangle\langle-1_{z}\rvert\bigr{)}+\bigl{(}1-w_{1}\bigr{)}\lvert 0_{z}\rangle\langle 0_{z}\rvert\,,$ (3)
where the $\lvert m_{z}\rangle$ are the eigenstates of $\hat{J}_{z}$. In a classical hadronization model at $t_{\mathrm{confine}}\ll t_{\mathrm{decay}}$ with intermediate state $p(\lambda)$, the probability of measuring the squared angular momentum $J_{e}^{2}=(\bm{e}\cdot\bm{J})^{2}$ along a unit vector $\bm{e}$ to be $x\in\{0,1\}$ reads
$P(x|\bm{e})=\int\mathrm{d}\lambda\,p(\lambda)\,P(x|\lambda,\bm{e})\,,\quad\langle J_{e}^{2}\rangle=P(1|\bm{e})$ (4)
and is subject to the constraint $\langle\bm{J}^{2}\rangle=2$ inherited from the interpolating partonic spin operator, i.e., $\langle J_{a}^{2}+J_{b}^{2}+J_{c}^{2}\rangle=2$ for any orthonormal triad $\{\bm{a},\bm{b},\bm{c}\}$. By Gleason’s theorem GLEASON (1957); Budroni et al. (2022), parity, and Lorentz covariance, $P(1|\bm{e})$ uniquely coincides with the single-parameter form of $\langle J_{e}^{2}\rangle$ as computed from Eq. (3) for any $\bm{e}$. All single-point spin measurements are thus determined by $P(1|\bm{z})=\langle J_{z}^{2}\rangle=w_{1}$, which again serves as a matching coefficient connecting the classical (or quantum-mechanical) model at $t_{\mathrm{confine}}$ to the quantum state at $t_{\mathrm{decay}}$.
Figure 2: QI characterization of the Falk-Peskin parameter $w_{1}$ for $Q\to\Sigma_{Q}^{*}X$ in terms of the KCBS inequality (blue), entanglement in a bipartite QM fragmentation model (orange), and classical correlation strength in a stochastic model (green). Experimental data are from Refs. Feindt (1995); Brandenburg et al. (1997),
In this context, noncontextuality of a hadronization model is the statement that in addition to Eq. (4), the model predicts certain joint probabilities
$P(x_{1},\dots,x_{N}|\bm{e}_{1},\dots,\bm{e}_{N})$ (5)
through a joint conditional measurement probability at given $\lambda$. Of particular interest are $\{\bm{e}_{i}\}$ arranged in connected Kochen-Specker (KS) graphs KOCHEN and SPECKER (1967) where $\bm{e}_{i}\perp\bm{e}_{j}$ for all edges in the graph. The first nontrivial case of a simple cyclic graph is the Klyachko-Can-Binicioglu-Shumovsky (KCBS) configuration =$\bigstar=\{\bm{e}_{i}\}$ Klyachko et al. (2008) of $N=5$ unit vectors $\bm{e}_{i}\perp\bm{e}_{i+1\!\!\!\mod 5}$ arranged around a symmetry axis $\bm{r}=\bm{z}$, i.e., they form a regular pentagram in the plane of fixed $\bm{r}\cdot\bm{e}_{i}=1/\sqrt[4]{5}$. Crucially, even though the additional measurements in Eq. (5) are counterfactual in our case, the existence of the joint distribution does leave an imprint on experimentally observable distributions. This is the statement of the KCBS inequality Klyachko et al. (2008), which in the general case constrains the sum of two-point correlations. Using $\langle\bm{J}^{2}\rangle=2$, one readily shows (also in the classical case) that Eq. (5) exists for $\bigstar$ if and only if the single-point marginal probabilities satisfy $\sum_{\bm{e}\in\bigstar}P(1|\bm{e})=\sum_{\bm{e}\in\bigstar}\langle J_{e}^{2}\rangle\geq 3$. Importantly, these latter single-point measurements are accessible experimentally and determined by $w_{1}$. We thus find a KCBS bound for the Falk-Peskin parameter, see Fig. 2,
$w_{1}\geq w_{1}^{\bigstar}\equiv\frac{1}{2}-\frac{1}{2\sqrt{5}}\approx 0.276393\,,$ (6)
which is satisfied if and only if Eq. (5) exists for $\bigstar$. We have verified that the constraint from $\bigstar$ with $N=5$ is the strongest constraint on $w_{1}$ from any simple $N$-cycle symmetric around $\bm{z}$. (In practice, we also choose $\bm{r}$ such that $\sum_{e\in\bigstar}\langle J_{e}^{2}\rangle$ is minimized.)
It is important to realize that classical models do not have to be noncontextual, cf. Eq. (7). Instead, the degree of contextuality is a property of the hadronization model that can be assessed by inspecting its dynamics and coupling to the final-state hadron spin. (The intuition of the KCBS-noncontextual limit being “more classical” is valuable nevertheless, cf. the case of a thermal system at high temperatures, where classical and quantum dynamics become indistinguishable and isotropic, $w_{1}\to 2/3>w_{1}^{\bigstar}$.) Conversely, the degree to which hadronization models can be coupled to hadron spin in a noncontextual way has fundamental limits. The tightest such constraint for $j=1$ arises from the Yu-Oh configuration Yu and Oh (2012), a KS graph with $N=13$ containing several cycles, whose associated state-independent noncontextuality inequality when combined with $\langle\bm{J}^{2}\rangle=2$ rules out the existence of Eq. (5) for this case altogether.
IV Classical correlation strength
To explore further interpretations of $w_{1}$, consider the “hidden” state of a classical hadronization model that simply consists of a unit vector $\bm{\lambda}$ in the hadron rest frame. Evaluating $P(1|\bm{z})$ in this model, we find
$w_{1}$ $=\frac{2}{3}+r\,\Delta_{2}\,,\quad\Delta_{2}\equiv\int\!\mathrm{d}c_{\theta}\,p(c_{\theta})\,\Bigl{(}c_{\theta}^{2}-\frac{1}{3}\Bigr{)}\,,$ (7)
where $c_{\theta}\equiv\bm{z}\cdot\bm{\lambda}$ and $r$ is a free parameter governing the strength of the coupling between $J_{e}^{2}$ and $\bm{\lambda}$ in the Lorentz-invariant conditional measurement probability $P(1|\bm{\lambda},\bm{e})=2/3+r[(\bm{e}\cdot\bm{\lambda})^{2}-1/3]$. Since $0\leq P(1|\bm{\lambda},\bm{e})\leq 1$, we have $-1\leq r\leq\tfrac{1}{2}$. If $\bm{\lambda}$ indicates e.g. the average orientation of classical diquark angular momenta near the heavy quark, one physically expects a positive coupling $r>0$ to the observed angular momentum, in which case $w_{1}\geq 1/2$ since $\Delta_{2}\geq-1/3$, exhibiting KCBS noncontextuality. However, nothing a priori forbids $r<0$. Since $\Delta_{2}\leq 2/3$, we can have $w_{1}\to 0$ in the extreme limit $r\to-1$ and $p(c_{\theta})\to\delta(\lvert c_{\theta}\rvert-1)/2$. This is of interest as an example of a classical model that achieves maximal KCBS violation, and thus must be contextual.
Another important insight from this limit is that $w_{1}$ measures the degree of classical correlation within the model. To make this explicit, consider a bipartite classical model consisting of two unit vectors $\{\bm{\lambda}_{1},\bm{\lambda}_{2}\}$ that couple to the hadron spin through $\bm{\lambda}=(\bm{\lambda}_{1}+\bm{\lambda}_{2})/\lvert\bm{\lambda}_{1}+\bm{\lambda}_{2}\rvert$. In this case we can prove a general bound $w_{1}\geq(1-|\mathcal{C}_{\lambda_{1}\lambda_{2}}|)/2$, see Fig. 2, where
$\mathcal{C}_{\lambda_{1}\lambda_{2}}\equiv\Big{\langle}\frac{2\lambda_{1z}\lambda_{2z}-\bm{\lambda}_{1}\cdot\bm{\lambda}_{2}\,(\lambda_{1z}^{2}+\lambda_{2z}^{2})}{1-(\bm{\lambda}_{1}\cdot\bm{\lambda}_{2})^{2}}\Big{\rangle}\leq 1\,,$ (8)
is a nonlinear correlation coefficient that vanishes e.g. if $\bm{\lambda}_{1,2}$ are independent, $p(\bm{\lambda}_{1},\bm{\lambda}_{2})=p(\bm{\lambda}_{1})\,p(\bm{\lambda}_{2})$. Alternatively, we can bound $w_{1}$ for $r<0$ by linear correlations, leading to
$w_{1}\geq\min\Bigl{\{}\langle\bm{\lambda}_{1T}\cdot\bm{\lambda}_{2T}\rangle,\,\frac{2}{3}\Bigr{\}}\,,$ (9)
where $\bm{\lambda}_{iT}$ are the components of $\bm{\lambda}_{i}$ transverse to $\bm{z}$.
V Bipartite QM models and factorization
We conclude our analysis of Eq. (3) by interpreting it in a simple QM model of hadronization where the hadron is formed from the heavy quark and two light spin-$1/2$ degrees of freedom, which we can e.g. think of as constituent quarks. In this case it is interesting to ask whether the state of the light system is entangled or separable when the hadron is formed. To assess this, we convert Eq. (3) to a two-qubit density matrix (noting that the singlet state, which would correspond to the $\Lambda_{Q}$ baryon, is not populated at this binding energy), and again evaluate the PPT criterion Horodecki et al. (1996). The result for the lowest eigenvalue of the partial transpose is shown in red in Fig. 2: As expected, an overlap of $\operatorname{Tr}[\rho\,\lvert 0_{z}\rangle\langle 0_{z}\rvert]=1-w_{1}>1/2$ with $\lvert 0_{z}\rangle$, which acquires an interpretation as a triplet Bell state in this model, implies entanglement (and eventually leads to KCBS violation Klyachko et al. (2008)).
Importantly, an observation of entanglement at $w_{1}<1/2$ would imply that no factorization can exist (in the technical sense of the word in high-energy physics) in terms of any spin-$1/2$ degrees of freedom at cross-section level. To see this, compare to the prototypical factorization of the hadronic tensor for the Drell-Yan process at leading-power,
$W^{\mu\nu}_{\mathrm{DY}}\sim\operatorname{tr}[\gamma^{\mu}\Phi_{q}\gamma^{\nu}\Phi_{\bar{q}}]\,,$ (10)
where $\Phi_{q,\bar{q}}$ are (anti)quark correlation functions of the (possibly polarized) incoming protons, $\operatorname{tr}$ denotes a trace over Dirac indices, and $\mu,\nu$ encode the polarization of the virtual photon. Eq. (10) is manifestly separable with respect to the product $\mathcal{H}_{q}\otimes\mathcal{H}_{\bar{q}}$ of the (anti)quark Hilbert spaces. Conversely, the neutral polarization state $\tfrac{1}{\sqrt{2}}\bigl{(}\lvert\uparrow_{q}\downarrow_{\bar{q}}\rangle+\lvert\downarrow_{q}\uparrow_{\bar{q}}\rangle\bigr{)}$, which is maximally entangled, cannot have large overlap with Eq. (10), and indeed it is well known Levelt and Mulders (1994); Boer (2009); Bacchetta et al. (2019); Vladimirov et al. (2022); Ebert et al. (2022); Gamberg et al. (2022); del Castillo et al. (2024); Gao et al. (2025) that neutral photon polarization requires an additional spin-$1$ degree of freedom (and thus subleading-power factorization). The above argument provides a novel perspective on this fact from QI theory.
VI TMD heavy-quark fragmentation
Figure 3: QI characterization of $Q\to\Sigma_{Q}^{*}X$ with a TMD measurement. The heatmap indicates the effective longitudinal polarization $w_{1}^{\prime}$. The bounds at $1/2$ and $w_{1}^{\bigstar}$ relevant for its QI interpretation are to be compared to Fig. 2.
It is experimentally possible Brandenburg et al. (1997) that inclusive fragmentation $Q\to\Sigma_{Q}^{*}X$ in fact produces an isotropic spin state with $w_{1}\approx 2/3$. It is interesting to ask, therefore, whether breaking the azimuthal symmetry by measuring a transverse momentum $\bm{k}_{T}\perp\bm{z}$ provides access to richer spin-$1$ dynamics. Here the transverse momentum-dependent (TMD) observable $\bm{k}_{T}\gtrsim\Lambda_{\mathrm{QCD}}$, may e.g. be (a) the recoil of the heavy hadron pair in $e^{+}e^{-}\to\Sigma_{Q}^{*}H_{\bar{Q}}X$ von Kuk et al. (2023); Dai et al. (2024); von Kuk et al. (2024); Lee et al. (2025), (b) the hadron transverse momentum relative to a jet containing it Kang et al. (2020); Arratia et al. (2020); Lee et al. (2025), or (c) the transverse momentum of an additional fragmentation pion Bacchetta et al. (2023); Lee et al. (2025). In any of these cases, the most general light angular momentum state compatible with the symmetries reads Lee et al. (2025)
$\rho$ $=\frac{\mathbb{1}}{3}-\tfrac{1}{2}\hat{J}_{y}\,w^{(1,1)}(k_{T})+\bigl{(}\tfrac{3}{2}\hat{J}_{z}^{2}-\mathbb{1}\bigr{)}\,w^{(2,0)}(k_{T})$ (11)
$+\bigl{(}\hat{J}_{x}\hat{J}_{z}+\hat{J}_{z}\hat{J}_{x}\bigr{)}\,w^{(2,1)}(k_{T})+\bigl{(}\hat{J}_{x}^{2}-\hat{J}_{y}^{2}\bigr{)}\,w^{(2,2)}(k_{T})\,,$
where $k_{T}=\lvert\bm{k}_{T}\rvert$, $\bm{x}=\bm{k}_{T}/k_{T}$, $\bm{y}=\bm{z}\times\bm{x}$, and the $w^{(N,n)}$ are the nonperturbative coefficients.
Our previous discussion elegantly carries over if we define
$w_{1}^{\prime}\equiv\langle J_{z^{\prime}}^{2}\rangle\equiv\min_{\bm{e}}\,\langle J_{e}^{2}\rangle=\min_{\bm{e}}\,\operatorname{Tr}\bigl{[}\rho\,\hat{J}_{e}^{2}\bigr{]}\,,$ (12)
i.e., by finding the axis $\bm{z^{\prime}}$ that maximizes the neutral polarization rate $1-w_{1}^{\prime}$. A rotation from $\bm{z}$ to $\bm{z}^{\prime}$ in fact diagonalizes $\rho$ with eigenvectors $\lvert m_{z^{\prime}}\rangle$, so we can think of $\bm{z^{\prime}}$ as a novel polarization axis that is spontaneously generated by the fragmentation dynamics. Here we can make immediate use of $\bm{z}^{\prime}$ by repeating our previous QI analysis for $w_{1}^{\prime}$, as shown in Fig. 3, where $w^{(2,0)}=0$ corresponds to $w_{1}=2/3$ after integrating over $\bm{k}_{T}$. The value of $w_{1}^{\prime}$ is driven by $w^{(2,1)}$ and $w^{(2,2)}$, and even for $w_{1}=2/3$ we find physical regions in their parameter space with nontrivial QI interpretations $w_{1}^{\prime}<1/2$ or $w_{1}^{\prime}<w_{1}^{\bigstar}$, respectively. (Increasing $w^{(1,1)}$ shrinks the physical region, but leaves $w_{1}^{\prime}$ unchanged.) Notably, their scaling for $k_{T}\gg\Lambda_{\mathrm{QCD}}$ is model-independently predicted to be $w^{(N,n)}(k_{T})\sim(\Lambda_{\mathrm{QCD}}/k_{T})^{n}$ Lee et al. (2025). Therefore if heavy-quark fragmentation indeed exhibits “quantum” behavior at $k_{T}\sim\Lambda_{\mathrm{QCD}}$ as in Fig. 3, there would be the exciting possibility to observe a quantum-to-classical transition as $k_{T}$ increases.
VII Conclusions and outlook
In this letter we have initiated the study of applying tools from quantum information (QI) theory to the open problem of hadronization to distinguish between entire classes of classical or quantum-mechanical hadronization models at once. In Eq. (2), we have presented an inequality for light dihadron fragmentation that, if violated in data, would rule out classical hadronization dynamics altogether. We stress that our goal was not to test quantum foundations themselves: Rather, we adapted tools from QI theory to characterize the dynamics of the blackbox Hamiltonian of hadronization, and appealed to effective field theory to separate it from other (very likely quantum) dynamics in the problem. We then adapted the concept of contextuality to the study of hadronization, showcased our ideas using the Falk-Peskin fragmentation parameter $w_{1}$ in heavy baryon production, and further interpreted $w_{1}$ in terms of classical correlation strength and QM entanglement, with implications for factorizability. Incorporating these insights will impose practical constraints on future hadronization modeling and enable cross-pollination with ongoing QI research.
The experimental status of $w_{1}$ is unclear at present, see Fig. 2. Combining $\Sigma_{b}^{*\pm}$ baryons, DELPHI measured Feindt (1995)
$w_{1}^{\mathrm{DELPHI}}=-0.36\pm 0.30_{\mathrm{stat}}\pm 0.30_{\mathrm{syst}}\,,$ (13)
which translates to a 95% CL limit of $w_{1}\leq 0.25$ after imposing positivity. On the other hand, by combining $\Sigma_{c}^{*++}$ and $\Sigma_{c}^{*0}$ baryons (that have the identical light valence content and relative rates), CLEO found Brandenburg et al. (1997),
$w_{1}^{\mathrm{CLEO}}=0.71\pm 0.13\,,$ (14)
close to the isotropic limit $w_{1}=2/3$. This disagreement has, to our knowledge, not been resolved Falk (1996). We therefore strongly encourage a new measurement of $w_{1}$ using state-of-the-art experimental techniques. We also encourage experimental tests of spin correlations in nearby dihadrons and of the TMD generalizations of $w_{1}$ that we introduced, all of which possess deep physical interpretations within the new QI framework of hadronization that we have presented here.
Acknowledgements.
Acknowledgments:
We would like to thank Jordi Tura Brugués, Herbi Dreiner, Jonas Helsen, Piet Mulders, Marieke Postma, Iain Stewart, and Jesse Thaler for fruitful discussion, as well as the members of the Nikhef LHCb Group and Mick Mulder in particular. We gratefully acknowledge the hospitality of the Nikhef Theory Group and the Erwin Schrödinger Institute. K.L. and Z.S. were supported by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
Supplemental material
In this supplemental material we collect explicit expressions for the differential distributions of decay products stemming from the self-analyzing decays of the polarized hadron systems discussed in the main text. These expressions involve the nonperturbative fragmentation coefficients relevant for the analysis in the main text as free parameters, and allow for their experimental determination by fitting them to the observed distributions.
VII.1 Fragmentation to nearby dihadrons
VII.1.1 Factorization and general structure
Consider the cross section for producing a dihadron pair with general spin $s_{1,2}$ and magnetic quantum numbers $-s_{1}\leq m_{1},m_{1}^{\prime}\leq s_{1}$ and $-s_{2}\leq m_{2},m_{2}^{\prime}\leq s_{2}$ with respect to the fragmentation axis $\mathbf{z}$. (In the collinear approximation, the two hadron rest frames only differ by a boost along the $\mathbf{z}$ direction, such that its orientation is common.) We decompose the cross section for an arbitrary spin state in terms of operators $\hat{\mathcal{O}}$ on the product Hilbert space as
$\frac{\mathrm{d}\sigma_{m_{1}m_{1}^{\prime},m_{2}m_{2}^{\prime}}}{\mathrm{d}z_{1}\,\mathrm{d}z_{2}\,\mathrm{d}R_{T}}=\sum_{\mathcal{O}}\frac{\mathrm{d}\sigma_{\mathcal{O}}}{\mathrm{d}z_{1}\,\mathrm{d}z_{2}\,\mathrm{d}R_{T}}\,\mathcal{O}_{m_{1}m_{1}^{\prime},m_{2}m_{2}^{\prime}}\,,$ (S1)
where $\mathcal{O}_{m_{1}m_{1}^{\prime},m_{2}m_{2}^{\prime}}=\langle m_{1}m_{2}\lvert\hat{\mathcal{O}}\rvert m_{1}^{\prime}m_{2}^{\prime}\rangle$ and $z_{1},z_{2}$ are the longitudinal momentum fractions of hadrons $h_{1}$ and $h_{2}$, respectively. Here, $R_{T}$ is the magnitude of the transverse component of the difference of hadron momenta relative to their sum. For pairs of spin-$1/2$ baryons we have
$\{\mathcal{O}\}=\Bigl{\{}\frac{\mathbb{1}}{4},\,\hat{S}_{1,L}\otimes\hat{S}_{2,L},\,\hat{\mathbf{S}}_{1,T}\otimes\hat{\mathbf{S}}_{2,T}\Bigr{\}}\,.$ (S2)
as in Eq. (1), where the coefficients $\mathrm{d}\sigma_{\mathcal{O}}$ are also denoted by $\mathrm{d}\sigma$ (without subscript), $\mathrm{d}\sigma_{LL}$, and $\mathrm{d}\sigma_{TT}$, respectively. They are given in terms of dihadron fragmentation functions $D^{\mathcal{O}}_{i\to h_{1}h_{2}}$ as Metz and Vossen (2016); Pitonyak et al. (2024); Rogers et al. (2025); Pitonyak et al. (2025); Huang et al. (2025)
$\frac{\mathrm{d}\sigma_{\mathcal{O}}}{\mathrm{d}z_{1}\,\mathrm{d}z_{2}\,\mathrm{d}R_{T}}=\sum_{i}\!\int\mathrm{d}y\,C_{i}(y,\mu)\,D^{\mathcal{O}}_{i\to h_{1}h_{2}}\left(\frac{z_{1}}{y},\frac{z_{2}}{y},R_{T},\mu\right)\,,$ (S3)
where $C_{i}$ are the usual single-inclusive partonic hard coefficients. (In the case of hadronic collisions, they also contain the parton distribution functions.) Additionally, $C_{i}$ can include fiducial acceptance cuts on the overall direction of the near-collimated dihadron system; cuts on individual hadron energies or momenta are special cases of the bins we consider below.
We are interested in the yield $N_{\mathcal{O}}$ of dihadron pairs produced in a specific spin state $\mathcal{O}$ and bin of $(z_{1},z_{2})$, which is proportional to
$\frac{N_{\mathcal{O}}}{\mathcal{L}}=\sigma_{\mathcal{O}}$ $\equiv\int\!\mathrm{d}z_{1}\,\mathrm{d}z_{2}\int_{0}^{R_{T}^{\rm{cut}}}\!\mathrm{d}R_{T}\,\frac{\mathrm{d}\sigma_{\mathcal{O}}}{\mathrm{d}z_{1}\,\mathrm{d}z_{2}\,\mathrm{d}R_{T}}\,\Theta(z_{1},z_{2})$
$=\sum_{i}\int\!\mathrm{d}y\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}\,C_{i}(y,\mu)\,\Theta(yx_{1},yx_{2})\,\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D^{\mathcal{O}}_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}\,,$ (S4)
where $\Theta(z_{1},z_{2})$ implements the given bin and the normalization factor $\mathcal{L}$ is the total luminosity. By placing an upper cutoff $R_{T}^{\mathrm{cut}}\sim\Lambda_{\mathrm{QCD}}$ on the relative transverse momentum, we ensure that spin correlations are genuinely sourced by the nonperturbative fragmentation dynamics. The total yield $N$ is given by the $N_{\mathcal{O}}$ with $\mathcal{O}=\mathbb{1}/\operatorname{Tr}[\mathbb{1}]$ the identity on the product Hilbert space. The normalized density matrix of the hadron pair then reads
$\rho=\frac{1}{N}\sum_{\mathcal{O}}N_{\mathcal{O}}\,\mathcal{O}\,.$ (S5)
For spin-$1/2$ baryons this results in Eq. (1) if we define the shorthands
$\tilde{D}_{LL}=\frac{\sigma_{LL}}{\sigma}=\frac{\sum_{i}\int\!\mathrm{d}y\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}\,C_{i}(y,\mu)\,\Theta(yx_{1},yx_{2})\,\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D^{LL}_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}{\sum_{i}\int\!\mathrm{d}y\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}C_{i}(y,\mu)\,\Theta(yx_{1},yx_{2})\,\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}\,,$ (S6)
and similarly for $\tilde{D}_{TT}$. We stress that even though the DiFFs enter in a convolution with the perturbative scattering coefficient in this form, the strength of the spin correlation is still uniquely sourced at the low scale, since the DGLAP evolution and the perturbative corrections are all common. In particular, using the leading-order perturbative result for the coefficient $C_{i}(y,\mu)\propto\delta(1-y)+\mathcal{O}(\alpha_{s})$, Eq. (S6) indeed reduces to a ratio of the (binned) DiFFs themselves.
We point out that while the expression for $\tilde{D}_{LL}$ in Eq. (S6) is clearly the most amenable to experimental tests of Eq. (2) in the near future, since it directly relates to the measured cross section at a given collider, we believe it would be valuable in the long run to test Eq. (2) directly at the level of the process and collider-independent ratio
$\tilde{D}_{LL,i}\bigl{(}x_{1},x_{2},R_{T}^{\mathrm{cut}}\bigr{)}\equiv\frac{\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D^{LL}_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}{\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}\,,$ (S7)
with $\tilde{D}_{TT,i}$ again defined in full analogy, where the right-hand side is evaluated at some low input scale $\mu\sim 2\,\mathrm{GeV}$. (In a more sophisticated setup, one may also define $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ as ratios of Mellin moments of the singlet and gluon or valence DiFFs to form exact renormalization-group invariants.) Eq. (2) equally well applies to $\tilde{D}_{LL,i}$ and $\tilde{D}_{TT,i}$ when defined as in Eq. (S7), where the normalized density matrix is now directly computed in terms of the renormalized fragmentation correlator for given hadron-hadron spin state at the low scale. Using Eq. (S7) in practice requires one to first determine the underlying DiFFs from a global fit to the available collider data for the yields $N_{\mathcal{O}}$ extracted from decay distributions in many different bins. Nevertheless, we believe this would be worth the effort it entails, since one would gain access to the degree of CHSH violation also as a function of e.g. the parton type.
VII.1.2 Decay distribution for nearby hyperon pairs
Differential distributions for the decay products of the hadrons are obtained by tracing Eq. (1) against the (conjugate) decay matrix elements for given hadron helicities $m_{1,2}$ ($m_{1,2}^{\prime}$), which are specific to each hadron and decay. Here we consider the parity-violating weak decay of hyperons, a standard example of a self-analyzing decay commonly used for reconstructing (transverse) hadron polarization, see e.g. Refs. Ablikim et al. (2019); Gong et al. (2022); Gamberg et al. (2021); Ablikim et al. (2022); Zhang and Wei (2023); Gao (2024); Wu et al. (2024b).
To fix conventions we begin with the decay of a single, individually polarized (anti)hyperon
$h_{1}=\Lambda\,\to p\,\pi^{-}\,,\qquad h_{2}=\bar{\Lambda}\,\to\bar{p}\,\pi^{+}\,.$ (S8)
The associated decay distributions are standard Zyla et al. (2020) and read
$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta_{p}\,\mathrm{d}\varphi_{p}}=\frac{1}{4\pi}\bigl{[}1+\alpha_{-}\,\bm{P}\cdot\bm{n}_{p}\bigr{]}\,,\qquad\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta_{\bar{p}}\,\mathrm{d}\varphi_{\bar{p}}}=\frac{1}{4\pi}\bigl{[}1+\alpha_{+}\,\bm{P}\cdot\bm{n}_{\bar{p}}\bigr{]}\,,$ (S9)
where $\bm{P}=2\langle\hat{\bm{S}}\rangle$ with $\lvert\bm{P}\rvert\leq 1$ is the Bloch vector of the (anti)hyperon, $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) is a unit vector pointing in the direction of the (anti)proton in the rest frame of the parent particle, $\theta_{p}$ and $\varphi_{p}$ ($\theta_{\bar{p}}$ and $\varphi_{\bar{p}}$) are the polar coordinates of $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) in the respective rest frame, and $\alpha_{\mp}$ are the hyperon decay parameters. With the above common sign for $\alpha_{\pm}$ for hyperon and antihyperon, one has opposite signs for $\alpha_{-}=0.732\pm 0.014$ and $\alpha_{+}=-0.758\pm 0.012$ Zyla et al. (2020) (and thus $\alpha_{-}\alpha_{+}<0$), as required by the approximate CP symmetry.
For the system in Eq. (1), each individual hyperon is indeed unpolarized on average, $\bm{P}=0$, such that the distributions $\mathrm{d}N/(\mathrm{d}\cos\theta_{p}\,\mathrm{d}\varphi_{p})$ and $\mathrm{d}N/(\mathrm{d}\cos\theta_{\bar{p}}\,\mathrm{d}\varphi_{\bar{p}})$ are flat. Nevertheless, the spin correlations encoded in Eq. (1) can be readily observed from the joint distribution of their decay products,
$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta_{p}\,\mathrm{d}\varphi_{p}\,\mathrm{d}\cos\theta_{\bar{p}}\,\mathrm{d}\varphi_{\bar{p}}}$ $=\frac{1}{(4\pi)^{2}}\bigl{[}1+\alpha_{-}\alpha_{+}\,\bm{n}_{p}\cdot\mathsf{T}\cdot\bm{n}_{\bar{p}}\bigr{]}$
$=\frac{1}{(4\pi)^{2}}\bigl{[}1+\alpha_{-}\alpha_{+}\,\tilde{D}_{LL}\,\cos\theta_{p}\,\cos\theta_{\bar{p}}+\alpha_{-}\alpha_{+}\,\tilde{D}_{TT}\,\sin\theta_{p}\,\sin\theta_{\bar{p}}\cos(\varphi_{p}-\varphi_{\bar{p}})\bigr{]}\,,$ (S10)
where the rank-two tensor $\mathsf{T}=(T_{jk})$ has components $T_{jk}/4=\langle S_{1,j}\otimes S_{2,k}\rangle$.
VII.2 Inclusive fragmentation to excited heavy baryons
The factorization theorem for the production of a boosted polarized heavy hadron $H$ containing a heavy quark of mass $m_{Q}$ at strict leading power in $\Lambda_{\mathrm{QCD}}/m_{Q}$ is simply a product Mele and Nason (1991); Jaffe and Randall (1994); Falk and Peskin (1994)
$\frac{\mathrm{d}\sigma_{HX}^{\mathcal{O}_{\ell}}}{\mathrm{d}^{3}\bm{P}_{H}}=\frac{\mathrm{d}\sigma_{QX}}{\mathrm{d}^{3}\bm{P}_{Q}}\,\chi_{H,\mathcal{O}_{\ell}}\Bigl{[}1+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m_{Q}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{m_{Q}}{\lvert\bm{P}_{H}\rvert}\Bigr{)}\Bigr{]}\,.$ (S11)
The matching coefficient $\mathrm{d}\sigma_{QX}$ is the differential cross section for producing a free unpolarized heavy quark at the same three-momentum in the center-of-mass frame of the collision, $\bm{P}_{Q}=\bm{P}_{H}$. The nonperturbative dynamics are contained in the fragmentation coefficient $\chi_{H,\mathcal{O}_{\ell}}$. Note that for hadron colliders, the second set of power corrections in Eq. (S11) scales as ${m_{Q}}/{\lvert\bm{P}_{T}\rvert}$ instead, where $\bm{P}_{T}$ is the transverse momentum relative to the beam axis. As before, $\mathcal{O}_{\ell}$ labels a spin operator encoding a generic spin state, in this case for the total angular momentum of the light constituents $\ell$ of $H$. The total probability $\chi_{H}$ for $Q$ to fragment into $H$, which is the coefficient $\chi_{H,\mathcal{O}}$ of $\mathcal{O}=\mathbb{1}_{\ell}/\operatorname{Tr}[\mathbb{1}_{\ell}]$, satisfies $\sum_{H}\chi_{H}=1$. Given the multiplicative form of Eq. (S11), cuts acting on the heavy hadron kinematics simply act on the heavy-quark production cross section. Thus, unlike the light dihadron case, no convolution is required to compute the total yields for each spin state. Instead, we have
$\rho=\frac{1}{N}\sum_{\mathcal{O}_{\ell}}N_{\mathcal{O}_{\ell}}\,\mathcal{O}_{\ell}=\frac{1}{\chi_{H}}\sum_{\mathcal{O}_{\ell}}\chi_{H,\mathcal{O}_{\ell}}\,\mathcal{O}_{\ell}\,,$ (S12)
i.e., the normalized spin density matrix of the light constituents in any given bin of the heavy-hadron kinematics is directly given by ratios of the nonperturbative coefficients. The normalized coefficients $\chi_{H,\mathcal{O}_{\ell}}/\chi_{H}$ for a light state with angular momentum $j=1$ or $j=3/2$ are in direct correspondence to the usual Falk-Peskin parameters $w_{j}$. Specifically, for the set of operators $\{\mathcal{O}_{\ell}\}=\{\frac{\mathbb{1}}{4},\,\hat{J}_{z}^{2}\}$ allowed by the symmetries in the case $j=1$, one readily converts the coefficients to the form involving $w_{1}$ in Eq. (3) using $\hat{J}_{z}^{2}=\mathbb{1}-\lvert 0_{z}\rangle\langle 0_{z}\rvert$.
Experimentally, the Falk-Peskin parameter $w_{1}$ can be measured from the pion distribution in the dominant decay $\Sigma_{Q}^{*}\to\Lambda_{Q}\pi$. Consistently working to the leading-order in the heavy-quark expansion, the decay amplitude is given by the Isgur-Wise transition matrix element Isgur and Wise (1991), which is proportional to
$Y_{1m}(\theta,\varphi)\big{\langle}1m\,\tfrac{1}{2}k\big{|}s\,h\big{\rangle}\,.$ (S13)
Here $Y_{jm}$ is a spherical harmonic, $\theta$ ($\varphi$) is the pion polar (azimuthal) angle relative to the fragmentation axis in the candidate $\Sigma_{Q}^{(*)}\to\Lambda_{Q}\pi$ rest frame, $s=1/2$ ($3/2$) and $h$ are the total spin and spin along the $\bm{z}$ axis of the $\Sigma_{Q}$ ($\Sigma_{Q}^{*}$) baryon, $k$ is the spin of the final-state $\Lambda_{Q}$ along the $\bm{z}$ axis, and $m=h-k$. We now dress Eqs. (3) and (S13) with the appropriate Clebsch-Gordan coefficients, average over the helicities of the unpolarized initial-state heavy-quark, and sum over the helicities of the final-state $\Lambda_{Q}$ (whose polarization is not reconstructed). This results in the following angular distributions for the decay products Falk and Peskin (1994),
$\frac{1}{N_{\Sigma_{Q}}}\frac{\mathrm{d}N_{\Sigma_{Q}}}{\mathrm{d}\cos\theta}=\frac{1}{2}\,,\qquad\frac{1}{N_{\Sigma_{Q}^{*}}}\frac{\mathrm{d}N_{\Sigma_{Q}^{*}}}{\mathrm{d}\cos\theta}$ $=\frac{1}{4}\Bigl{[}1+3\cos^{2}\theta-\frac{9}{2}\omega_{1}\Bigl{(}\cos^{2}\theta-\frac{1}{3}\Bigr{)}\Bigr{]}$
$=\frac{1}{2}\biggl{\{}1-\frac{3}{8}\bigl{[}1+3\cos(2\theta)\bigr{]}\Bigl{(}w_{1}-\frac{2}{3}\Bigr{)}\biggr{\}}\,.$ (S14)
Here we have performed the integral over the $\varphi$ dependence, which as expected is flat since no azimuthal direction is preferred. Here we assumed that the $\Sigma_{Q}$ and $\Sigma_{Q}^{*}$ form well-separated resonances $\Gamma_{\Sigma_{Q}},\Gamma_{\Sigma_{Q}^{*}}\ll m_{\Sigma_{Q}^{*}}-m_{\Sigma_{Q}}$ relative to the mass splitting in the heavy-quark spin symmetry doublet, which is reasonably well satisfied Aaltonen et al. (2012); Aaij et al. (2019). If needed, the decay distributions can readily be generalized to account for their interference Falk and Peskin (1994), and still can be expressed purely in terms of $w_{1}$ in that case. We note that the continuum $Q\to\Lambda_{Q}\pi$ background to Eq. (VII.2) is in fact described by a heavy-light DiFF that is of interest of its own, with the pion polar angle tracking its residual momentum fraction $z_{\pi}$ in this case. Importantly, however, no feature appears around $m_{\Lambda_{Q}\pi}\approx m_{\Sigma_{Q}^{*}}$ in the double-differential $(z_{\pi},m_{\Lambda_{Q}\pi})$ distribution in the background case. We therefore expect that a standard sideband subtraction will remain viable to remove the background also bin by bin in $\theta$. (Signal-background interference is likewise suppressed by the narrow widths of the resonances.) Finally, we stress that care must be taken to avoid biasing the angular distribution in Eq. (VII.2) when applying acceptance cuts on the pion transverse momentum with respect to the beam axis, as typically done in spectroscopic analyses Aaltonen et al. (2012); Aaij et al. (2019) to enrich the sample with signal events and suppress the $\Lambda_{Q}\pi$ continuum background. An attractive way to assess systematic uncertainties from correcting for the acceptance (as well as from subleading $1/m_{Q}$ corrections on the theory side) is offered by the corresponding angular distribution for $\Sigma_{Q}$ baryons, which is flat in $\cos\theta$ for unpolarized collisions at leading order in the heavy-quark expansion.
VII.3 Transverse-momentum dependent fragmentation to excited heavy baryons
The precise relation of the coefficients $w^{(N,n)}$ in Eq. (11) to underlying renormalized nonperturbative matrix elements differs, depending on the case considered: In case (c) given in the main text, corresponding to heavy-light dihadron fragmentation, the factorization takes the same multiplicative form as in Eq. (S11). For cases (a) and (b) the matrix elements enter through convolutions with other TMD matrix elements encoding the contributions of soft dynamics or other collinear sectors, none of which however affect the boosted heavy-hadron spin dynamics or the symmetry analysis leading to Eq. (11). Importantly, the resulting decay pion distribution for unpolarized collisions takes a common general form for any of these cases in terms of the cross section-level spin density in Eq. (11), and evaluates to
$\frac{1}{\mathrm{d}N_{\Sigma_{Q}}/\mathrm{d}k_{T}}\frac{\mathrm{d}N_{\Sigma_{Q}}}{\mathrm{d}k_{T}\,\mathrm{d}\cos\theta\,\mathrm{d}\varphi}$ $=\frac{1}{4\pi}\,,$ (S15)
$\frac{1}{\mathrm{d}N_{\Sigma_{Q}^{*}}/\mathrm{d}k_{T}}\frac{\mathrm{d}N_{\Sigma_{Q}^{*}}}{\mathrm{d}k_{T}\,\mathrm{d}\cos\theta\,\mathrm{d}\varphi}$ $=\frac{1}{4\pi}\biggl{\{}1-\frac{3}{8}\bigl{[}1+3\cos(2\theta)\bigr{]}\,w^{(2,0)}-3\cos\theta\,\sin\theta\,\cos\varphi\,w^{(2,1)}-\frac{3}{2}\sin^{2}\theta\,\cos(2\varphi)\,w^{(2,2)}\biggr{\}}\,,$
where $k_{T}$ is the magnitude of the transverse momentum $\bm{k}_{T}\perp\bm{z}$, reconstructed according to the respective experimental scenario. Importantly, $\varphi$ in this case is the azimuthal angle of the pion relative to the azimuth of $\bm{k}_{T}$, which provides an azimuthal reference direction $\bm{x}=\bm{k}_{T}/k_{T}$ in this case, as discussed in the main text.

### v2

@@
 Calculating this nonperturbative, strongly coupled phenomenon called hadronization from first principles using the Lagrangian of Quantum Chromodynamics (QCD) remains an open problem, restricting us to phenomenological models.
-Perhaps surprisingly, and despite the underlying dynamics of quark and gluons clearly being quantum mechanical (QM) in general, classical stochastic models of hadronization Field and Feynman (1978); Field (1989); Andersson et al.
+Perhaps surprisingly, and despite the underlying dynamics of quarks and gluons clearly being quantum mechanical (QM) in general, classical stochastic models of hadronization Field and Feynman (1978); Field (1989); Andersson et al.
 (1983); Webber (1984) as e.g.
@@
 Our proposal is part of a larger fruitful effort to bring QI theory to bear on questions in particle physics Tornqvist (1981); Ablikim et al.
+(2019); Neill and Waalewijn (2019); Afik and de Nova (2021); Takubo et al.
 (2021); Severi et al.
-(2022); Aoude et al.
-(2022); Afik and de Nova (2023); Aad et al.
+(2022); Afik and de Nova (2022); Aoude et al.
+(2022); Afik and de Nova (2023); Dong et al.
+(2024); Morales (2023); Aad et al.
 (2024); Barr et al.
@@
 (2024b); Fabbrichesi et al.
-(2024); Demina and Landi (2025); Ruzi et al.
+(2024); Demina and Landi (2025); Lo Chiatto (2024); Ruzi et al.
 (2024); Morales and Szynkman (2024); Grossi et al.
@@
 (2025a); Cheng and Yan (2025); Fabbrichesi et al.
-(2025b, c), but distinct in that most of the existing work, if not attempting to explain the Standard Model (SM) Mulders (2018); Thaler and Trifinopoulos (2024), aims to probe quantum foundations either at the highest energy scales or in exclusive hadronic processes, essentially asserting the quantum nature of the perturbative SM Lagrangian or hadronic form factors, respectively.
+(2025b, c), but distinct in that most of the existing work, if not attempting to explain the Standard Model (SM) Cervera-Lierta et al.
+(2017); Mulders (2018); Thaler and Trifinopoulos (2025), aims to probe quantum foundations either at the highest energy scales or in exclusive hadronic processes, essentially asserting the quantum nature of the perturbative SM Lagrangian or hadronic form factors, respectively.
 Hadronization then either serves as an analyzer for the perturbative density matrix Aidala and Rogers (2021); Afik et al.
@@
 (2024); Li et al.
-(2024); Lee et al.
-(2024); Bauer (2025).
+(2024a); Lee et al.
+(2025a); Bauer (2025).
 II Classical hadronization models as “hidden” variables
@@
 (1983); Webber (1984), one can print out and bin the intermediate state to record $p(\lambda)$.
-Just as in Bell’s concise proof Bell and Aspect (2004) of the Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al.
-(1969), suitable correlation observables $\mathcal{C}$ will have an expectation value $\langle\mathcal{C}\rangle_{p(\lambda)}\leq\mathcal{C}_{\mathrm{max}}$ that is bounded when computed in any classical hadronization model, following basic properties of the integral measure in any number of classical phase-space dimensions, while QM expectation values can exceed the bound.
+One then shows that for large classes of such models, suitable correlation observables $\mathcal{C}$ have a bounded expectation value $\langle\mathcal{C}\rangle_{p(\lambda)}\leq\mathcal{C}_{\mathrm{max}}$, using basic properties of the integral measure in any number of classical phase-space dimensions, while QM expectation values can exceed the bound.
 Note that $p(\lambda)$ and the conditional measurement probabilities of classical hadronization models must obey the spacetime symmetries of QCD (Lorentz covariance and parity), a constraint that is typically absent in tests of quantum foundations, but which we will make productive use of in this letter.
@@
 (In the last term, a contraction between vector indices is understood.) Eq.
-(1) is the polarization state of the system at long distances of order the hadron lifetimes, $t_{\mathrm{decay}}\sim 1/\Gamma_{h}$, where a QM description of the system is appropriate for the exclusive hadronic decays.111 Decay matrix elements and differential distributions are collected in the supplemental material.
+(1) is the polarization state of the system at long distances of order the hadron lifetimes, $t_{\mathrm{decay}}\sim 1/\Gamma_{h}$, where a QM description of the system is appropriate for the exclusive hadronic decays sup .
 By contrast, the nonperturbative hadronization physics are encoded in normalized dihadron fragmentation functions (DiFFs) Metz and Vossen (2016); Pitonyak et al.
@@
 (1996), the hadrons are entangled if and only if $\lvert\tilde{D}_{TT}\rvert>(\tilde{D}_{LL}-1)/2$.
-Most interestingly, we can evaluate the optimal Horodecki et al.
-(1995) CHSH inequality Clauser et al.
-(1969) for Eq.
-(1) in terms of the DiFFs,
-$\tilde{D}_{TT}^{2}+\max\{\tilde{D}_{LL}^{2},\tilde{D}_{TT}^{2}\}\stackrel{{\scriptstyle p(\lambda)}}{{\leq}}1\,.$ (2)
+Most interestingly, we can consider the class of fully classical hadronization models where the (spin) properties of individual final-state hadrons are separately determined by a common underlying classical state $\lambda$,
+$\langle S_{1,k}\otimes S_{2,\ell}\rangle=\int\!\mathrm{d}\lambda\,p(\lambda)\,S_{1,k}(\lambda)\,S_{2,\ell}(\lambda)\,.$ (2)
+For these models, the following optimal Horodecki et al. (1995) Clauser-Horne-Shimony-Holt (CHSH) inequality Clauser et al. (1969) holds,
+[…]
 This is the key result for this example process, illustrating our proposal: If Eq.
-(2) is found to be violated in data, the observation cannot be explained by any classical hadronization model.
-The proof exactly follows Ref.
-Bell and Aspect (2004), with one crucial difference: In our case, the “locality” assumption that measurements on $h_{1}$ are independent of the reconstruction of $h_{2}$ is guaranteed by color confinement, rather than by hypothesizing, like Bell, a local hidden-variable theory (LHVT) of Nature.
+(3) is found to be violated in data, the observation cannot be explained by any fully classical hadronization model of the form Eq.
+(2).111 By comparison, the observed Bose-Einstein enhancement of nearby identical hadrons Andersson and Hofmann (1986); Artru and Bowler (1988); Acton et al.
+(1993) simply follows from (anti)symmetrizing the wave function in momentum space.
+Doing so on top of a classical model leaves the normalized Eq.
+(1) unaffected, making Eq.
+(3) a deeper test of genuine QM dynamics.
+The derivation exactly follows Bell’s concise proof of the CHSH inequality in Ref.
+Bell and Aspect (2004), with one crucial difference: In our case, the “locality” assumption in Eq.
+(2) is a property that can be checked for any given hadronization model (without computation) by inspecting its dynamical rules and coupling to hadron spins, and not a property of a Bell-type local hidden-variable theory (LHVT) of Nature.
 The CHSH test we propose here therefore does not test quantum nonlocality in the usual way, but instead uses the identical mathematics to test whether the blackbox Hamiltonian of hadronization admits a classical description.222 Another assumption is that Gedanken measurements $\langle(\bm{e}_{\alpha}\cdot\bm{S}_{1})\otimes(\bm{S}_{2}\cdot\bm{e}_{\beta})\rangle$ at Bell angles $\alpha,\beta$ are indeed predicted to evaluate to $\bm{e}_{\alpha}\cdot(\mathsf{T}/4)\cdot\bm{e}_{\beta}$, where $\mathsf{T}=(T_{k\ell})$.
@@
 (1970); Abel et al.
-(1992) (nota bene constructed to argue against the viability of testing LHVTs of Nature at colliders) do not apply.
+(1992); Li et al.
+(2024b) (nota bene constructed to argue against the viability of testing LHVTs of Nature at colliders) do not apply.
 Figure 1: QI characterization of parton fragmentation to two nearby polarized hadrons.
@@
 For $\tilde{D}_{LL}=-1$ and $\tilde{D}_{TT}=-1$ ($+1$), the hadrons are produced in a pure singlet (triplet) Bell state.
+For the experimental data, see Ref.
+STAR Collaboration (2025); sup .
 III Quantum contextuality and local spin-1 systems
@@
 This is the statement of the KCBS inequality Klyachko et al.
-(2008), which in the general case constrains the sum of two-point correlations.
+(2008), which involves two-point correlations.
 Using $\langle\bm{J}^{2}\rangle=2$, one readily shows (also in the classical case) that Eq.
-(5) exists for $\bigstar$ if and only if the single-point marginal probabilities satisfy $\sum_{\bm{e}\in\bigstar}P(1|\bm{e})=\sum_{\bm{e}\in\bigstar}\langle J_{e}^{2}\rangle\geq 3$.
+(6) exists for $\bigstar$ if and only if the single-point marginal probabilities satisfy $\sum_{\bm{e}\in\bigstar}\langle J_{e}^{2}\rangle\geq 3$.
 Importantly, these latter single-point measurements are accessible experimentally and determined by $w_{1}$.
@@
 (8).
-Instead, the degree of contextuality is a property of the hadronization model that can be assessed by inspecting its dynamics and coupling to the final-state hadron spin.
+Instead, the degree of contextuality, like Eq.
+(2), is a property of the hadronization model that can be assessed by inspecting its dynamics and coupling to the final-state hadron spin.
 (The intuition of the KCBS-noncontextual limit being “more classical” is valuable nevertheless, cf.
@@
 V Bipartite QM models and factorization
-We conclude our analysis of Eq.
-(3) by interpreting it in a simple QM model of hadronization where the hadron is formed from the heavy quark and two light spin-$1/2$ degrees of freedom, which we can e.g.
-think of as constituent quarks.
-In this case it is interesting to ask whether the state of the light system is entangled or separable when the hadron is formed.
+Lastly, we interpret Eq.
+(4) in a simple QM model of hadronization where the hadron is formed from the heavy quark and two light spin-$1/2$ degrees of freedom, e.g.
+constituent quarks.
+Here it is interesting to ask whether the light system is entangled or separable when the hadron is formed.
 To assess this, we convert Eq.
@@
 Eq.
-(10) is manifestly separable with respect to the product $\mathcal{H}_{q}\otimes\mathcal{H}_{\bar{q}}$ of the (anti)quark Hilbert spaces.
+(11) is manifestly separable with respect to the product $\mathcal{H}_{q}\otimes\mathcal{H}_{\bar{q}}$ of the (anti)quark Hilbert spaces Boer et al.
+(2005).
 Conversely, the neutral polarization state $\tfrac{1}{\sqrt{2}}\bigl{(}\lvert\uparrow_{q}\downarrow_{\bar{q}}\rangle+\lvert\downarrow_{q}\uparrow_{\bar{q}}\rangle\bigr{)}$, which is maximally entangled, cannot have large overlap with Eq.
@@
 (2024); Lee et al.
-(2025), (b) the hadron transverse momentum relative to a jet containing it Kang et al.
+(2025b), (b) the hadron transverse momentum relative to a jet containing it Kang et al.
 (2020); Arratia et al.
 (2020); Lee et al.
-(2025), or (c) the transverse momentum of an additional fragmentation pion Bacchetta et al.
+(2025b), or (c) the transverse momentum of an additional fragmentation pion Bacchetta et al.
 (2023); Lee et al.
-(2025).
+(2025b).
 In any of these cases, the most general light angular momentum state compatible with the symmetries reads Lee et al.
-(2025)
+(2025b)
 $\rho$ $=\frac{\mathbb{1}}{3}-\tfrac{1}{2}\hat{J}_{y}\,w^{(1,1)}(k_{T})+\bigl{(}\tfrac{3}{2}\hat{J}_{z}^{2}-\mathbb{1}\bigr{)}\,w^{(2,0)}(k_{T})$ (12)
@@
 (Increasing $w^{(1,1)}$ shrinks the physical region, but leaves $w_{1}^{\prime}$ unchanged.) Notably, their scaling for $k_{T}\gg\Lambda_{\mathrm{QCD}}$ is model-independently predicted to be $w^{(N,n)}(k_{T})\sim(\Lambda_{\mathrm{QCD}}/k_{T})^{n}$ Lee et al.
-(2025).
+(2025b).
 Therefore if heavy-quark fragmentation indeed exhibits “quantum” behavior at $k_{T}\sim\Lambda_{\mathrm{QCD}}$ as in Fig.
@@
 We therefore strongly encourage a new measurement of $w_{1}$ using state-of-the-art experimental techniques.
-We also encourage experimental tests of spin correlations in nearby dihadrons and of the TMD generalizations of $w_{1}$ that we introduced, all of which possess deep physical interpretations within the new QI framework of hadronization that we have presented here.
+We also encourage experimental tests of spin correlations in nearby dihadrons STAR Collaboration (2025) and of the TMD generalizations of $w_{1}$ that we introduced, all of which possess deep physical interpretations within the new QI framework of hadronization that we have presented here.
 Acknowledgements.
 Acknowledgments:
-We would like to thank Jordi Tura Brugués, Herbi Dreiner, Jonas Helsen, Piet Mulders, Marieke Postma, Iain Stewart, and Jesse Thaler for fruitful discussion, as well as the members of the Nikhef LHCb Group and Mick Mulder in particular.
-We gratefully acknowledge the hospitality of the Nikhef Theory Group and the Erwin Schrödinger Institute.
+We would like to thank Daniël Boer, Jordi Tura Brugués, Herbi Dreiner, Jonas Helsen, Piet Mulders, Duff Neill, Marieke Postma, Iain Stewart, and Jesse Thaler for fruitful discussion, as well as the members of the Nikhef LHCb Group and Mick Mulder in particular.
+The authors gratefully acknowledge the hospitality of the Nikhef Theory Group, the Institute for Theoretical Physics at the University of Amsterdam, and the Erwin Schrödinger Institute.
 K.L.
@@
 as in Eq.
-(1), where the coefficients $\mathrm{d}\sigma_{\mathcal{O}}$ are also denoted by $\mathrm{d}\sigma$ (without subscript), $\mathrm{d}\sigma_{LL}$, and $\mathrm{d}\sigma_{TT}$, respectively.
+(1), where the coefficients $\mathrm{d}\sigma_{\mathcal{O}}$ in turn are also denoted by $\mathrm{d}\sigma$ (without subscript), $\mathrm{d}\sigma_{LL}$, and $\mathrm{d}\sigma_{TT}$.
 They are given in terms of dihadron fragmentation functions $D^{\mathcal{O}}_{i\to h_{1}h_{2}}$ as Metz and Vossen (2016); Pitonyak et al.
@@
 $\rho=\frac{1}{N}\sum_{\mathcal{O}}N_{\mathcal{O}}\,\mathcal{O}\,.$ (S5)
-For spin-$1/2$ baryons this results in Eq. (1) if we define the shorthands
+For spin-$1/2$ baryons this results in Eq. (1) if we define the shorthand
 $\tilde{D}_{LL}=\frac{\sigma_{LL}}{\sigma}=\frac{\sum_{i}\int\!\mathrm{d}y\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}\,C_{i}(y,\mu)\,\Theta(yx_{1},yx_{2})\,\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D^{LL}_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}{\sum_{i}\int\!\mathrm{d}y\,\mathrm{d}x_{1}\,\mathrm{d}x_{2}C_{i}(y,\mu)\,\Theta(yx_{1},yx_{2})\,\int_{0}^{R_{T}^{\mathrm{cut}}}\!\mathrm{d}R_{T}\,D_{i\to h_{1}h_{2}}\bigl{(}x_{1},x_{2},R_{T},\mu\bigr{)}}\,,$ (S6)
@@
 (3) equally well applies to $\tilde{D}_{LL,i}$ and $\tilde{D}_{TT,i}$ when defined as in Eq.
-(S7), where the normalized density matrix is now directly computed in terms of the renormalized fragmentation correlator for given hadron-hadron spin state at the low scale.
+(S7), where the normalized density matrix is now directly computed in terms of the renormalized dihadron fragmentation correlator at the low scale.
 Using Eq.
@@
 where $\bm{P}=2\langle\hat{\bm{S}}\rangle$ with $\lvert\bm{P}\rvert\leq 1$ is the Bloch vector of the (anti)hyperon, $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) is a unit vector pointing in the direction of the (anti)proton in the rest frame of the parent particle, $\theta_{p}$ and $\varphi_{p}$ ($\theta_{\bar{p}}$ and $\varphi_{\bar{p}}$) are the polar coordinates of $\bm{n_{p}}$ ($\bm{n_{\bar{p}}}$) in the respective rest frame, and $\alpha_{\mp}$ are the hyperon decay parameters.
-With the above common sign for $\alpha_{\pm}$ for hyperon and antihyperon, one has opposite signs for $\alpha_{-}=0.732\pm 0.014$ and $\alpha_{+}=-0.758\pm 0.012$ Zyla et al.
+With the above common sign in front of $\alpha_{\pm}$ for hyperon and antihyperon, one has opposite signs for $\alpha_{-}=0.732\pm 0.014$ and $\alpha_{+}=-0.758\pm 0.012$ Zyla et al.
 (2020) (and thus $\alpha_{-}\alpha_{+}<0$), as required by the approximate CP symmetry.
@@
 $=\frac{1}{(4\pi)^{2}}\bigl{[}1+\alpha_{-}\alpha_{+}\,\tilde{D}_{LL}\,\cos\theta_{p}\,\cos\theta_{\bar{p}}+\alpha_{-}\alpha_{+}\,\tilde{D}_{TT}\,\sin\theta_{p}\,\sin\theta_{\bar{p}}\cos(\varphi_{p}-\varphi_{\bar{p}})\bigr{]}\,,$ (S10)
-where the rank-two tensor $\mathsf{T}=(T_{jk})$ has components $T_{jk}/4=\langle S_{1,j}\otimes S_{2,k}\rangle$.
+where the rank-two tensor $\mathsf{T}=(T_{k\ell})=\operatorname{diag}(\tilde{D}_{TT},\tilde{D}_{TT},\tilde{D}_{LL})$ has components $T_{k\ell}/4=\langle S_{1,k}\otimes S_{2,\ell}\rangle$.
+We note that Eq.
+(VII.1.2) takes the same form for the decay products of same-sign hyperon pairs (or other exclusive baryonic decays that follow Eq.
+(S9)) after suitably replacing the decay constants.
+Comparison to Ref. STAR Collaboration (2025):
+Very recently, and after the first preprint version of this manuscript was made public, the STAR collaboration published a measurement STAR Collaboration (2025) of the differential distribution
+$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta^{*}}=\frac{1}{2}\bigl{[}1+\alpha_{-}\alpha_{+}\,P_{\Lambda\bar{\Lambda}}\,\cos\theta^{*}\bigr{]}\,,$ (S11)
+where in our notation $\cos\theta^{*}=\bm{n}_{p}\cdot\bm{n}_{\bar{p}}$ and the coefficient $P_{\Lambda\bar{\Lambda}}$ encodes the strength of the angular correlation between the decay (anti)protons, and thus the strength of the spin correlation between the parent (anti)hyperons.
+Projecting our Eq.
+(VII.1.2) onto the distribution differential in $\cos\theta^{*}$, we find the relation
+$3P_{\Lambda\bar{\Lambda}}=\operatorname{Tr}\mathsf{T}=2\tilde{D}_{TT}+\tilde{D}_{LL}\,.$ (S12)
+The measurement of $P_{\Lambda\bar{\Lambda}}$ in Ref.
+STAR Collaboration (2025), as well as those for same-sign hyperon pairs obtained in Ref.
+STAR Collaboration (2025), therefore experimentally constrain a very specific linear combination of dihadron fragmentation functions.
+Ref.
+STAR Collaboration (2025) found that for nearby dihadrons, which as discussed in the main text and the above supplemental material is the most interesting kinematic region, the spin correlation $P_{\Lambda\bar{\Lambda}}=0.181\pm 0.035_{\mathrm{stat}}\pm 0.022_{\mathrm{sys}}$ was near its maximal value of $1/3$ (with some loss expected due to feed-down contributions from higher resonances), while $P_{\Lambda\Lambda}$ and $P_{\bar{\Lambda}\bar{\Lambda}}$ were still compatible with zero.
+These experimental constraints are indicated in Fig.
+1 as of the second preprint version of this manuscript.
+Notably, either of the experimental results is still compatible with entanglement or even CHSH violation in the upper left corner, since those depend on the precise breakdown between $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ and in particular the sign of $\tilde{D}_{LL}$.
+We stress that two caveats apply when interpreting the measurement of Ref.
+STAR Collaboration (2025) in this way; specifically, (a) further work is needed to ascertain the validity of the factorization in terms of fragmentation functions, which might require a more stringent cut on the $p_{T}$ of the hyperon candidates, and (b) the equivalence between the respective reference frames and angles should be carefully checked.
+The analysis in Ref.
+STAR Collaboration (2025) was in part motivated by theory considerations that postulate a rotationally invariant distribution of decay products, as e.g.
+appropriate for exclusive $\Lambda\bar{\Lambda}$ production, in which case $\cos\theta^{*}$ is the only angular observable of interest.
+As we have discussed, however, the fragmentation process necessarily retains a dependence on the fragmentation axis $\hat{z}$, i.e., the orientation relative to the hard scattering process that produces the parent parton, and whose hard scattering products balance the parton’s color charge.
+For this reason the polarization tensor T need not be proportional to the identity, and indeed in general has two nontrivial degrees of freedom ($\tilde{D}_{LL}$ and $\tilde{D}_{TT}$) according to our symmetry analysis that led to Eq.
+(1).
+We therefore strongly encourage an extension of the exciting measurement reported in Ref.
+STAR Collaboration (2025) to the more differential angular distribution in Eq.
+(VII.1.2), which would break the degeneracy between $\tilde{D}_{LL}$ and $\tilde{D}_{TT}$ and hold the potential of a definitive interpretation in terms of QI constraints on possible hadronization models.
+Alternatively, the degeneracy could at this point also be broken by a simpler, direct measurement of $\tilde{D}_{LL}$ via the $\cos\theta_{p}\cos\theta_{\bar{p}}$ correlation in Eq.
+(VII.1.2) in a way that is inclusive over $\phi_{p,\bar{p}}$, i.e.,
+$\frac{1}{N}\frac{\mathrm{d}N}{\mathrm{d}\cos\theta_{p}\,\mathrm{d}\cos\theta_{\bar{p}}}$ $=\frac{1}{4}\bigl{[}1+\alpha_{-}\alpha_{+}\,\tilde{D}_{LL}\,\cos\theta_{p}\,\cos\theta_{\bar{p}}\bigr{]}\,.$ (S13)
 VII.2 Inclusive fragmentation to excited heavy baryons
-The factorization theorem for the production of a boosted polarized heavy hadron $H$ containing a heavy quark of mass $m_{Q}$ at strict leading power in $\Lambda_{\mathrm{QCD}}/m_{Q}$ is simply a product Mele and Nason (1991); Jaffe and Randall (1994); Falk and Peskin (1994)
+The factorization theorem for the production of a boosted polarized heavy hadron $H$ containing a heavy quark of mass $m_{Q}$ at strict leading power in $\Lambda_{\mathrm{QCD}}/m_{Q}$ is simply a product Mele and Nason (1991); Jaffe and Randall (1994); Falk and Peskin (1994),
 $\frac{\mathrm{d}\sigma_{HX}^{\mathcal{O}_{\ell}}}{\mathrm{d}^{3}\bm{P}_{H}}=\frac{\mathrm{d}\sigma_{QX}}{\mathrm{d}^{3}\bm{P}_{Q}}\,\chi_{H,\mathcal{O}_{\ell}}\Bigl{[}1+\mathcal{O}\Bigl{(}\frac{\Lambda_{\mathrm{QCD}}}{m_{Q}}\Bigr{)}+\mathcal{O}\Bigl{(}\frac{m_{Q}}{\lvert\bm{P}_{H}\rvert}\Bigr{)}\Bigr{]}\,.$ (S14)
@@
 The normalized coefficients $\chi_{H,\mathcal{O}_{\ell}}/\chi_{H}$ for a light state with angular momentum $j=1$ or $j=3/2$ are in direct correspondence to the usual Falk-Peskin parameters $w_{j}$.
-Specifically, for the set of operators $\{\mathcal{O}_{\ell}\}=\{\frac{\mathbb{1}}{4},\,\hat{J}_{z}^{2}\}$ allowed by the symmetries in the case $j=1$, one readily converts the coefficients to the form involving $w_{1}$ in Eq.
+Specifically, for the set of operators $\{\mathcal{O}_{\ell}\}=\{\mathbb{1}/4,\,\hat{J}_{z}^{2}\}$ allowed by the symmetries in the case $j=1$, one readily converts the coefficients to the form involving $w_{1}$ in Eq.
 (4) using $\hat{J}_{z}^{2}=\mathbb{1}-\lvert 0_{z}\rangle\langle 0_{z}\rvert$.
@@
 Here $Y_{jm}$ is a spherical harmonic, $\theta$ ($\varphi$) is the pion polar (azimuthal) angle relative to the fragmentation axis in the candidate $\Sigma_{Q}^{(*)}\to\Lambda_{Q}\pi$ rest frame, $s=1/2$ ($3/2$) and $h$ are the total spin and spin along the $\bm{z}$ axis of the $\Sigma_{Q}$ ($\Sigma_{Q}^{*}$) baryon, $k$ is the spin of the final-state $\Lambda_{Q}$ along the $\bm{z}$ axis, and $m=h-k$.
+We recall the following, very simple form of the fragmentation axis $\bm{z}$ when expressed in terms of the beam three-momenta $\bm{P}_{a,b}$ in the candidate rest frame,
+$\bm{z}=-\frac{\bm{P_{a}}+\bm{P_{b}}}{\lvert\bm{P_{a}}+\bm{P_{b}}\rvert}\,,$ (S17)
+where the beams can be protons or leptons.
 We now dress Eqs.
@@
 Here we have performed the integral over the $\varphi$ dependence, which as expected is flat since no azimuthal direction is preferred.
-Here we assumed that the $\Sigma_{Q}$ and $\Sigma_{Q}^{*}$ form well-separated resonances $\Gamma_{\Sigma_{Q}},\Gamma_{\Sigma_{Q}^{*}}\ll m_{\Sigma_{Q}^{*}}-m_{\Sigma_{Q}}$ relative to the mass splitting in the heavy-quark spin symmetry doublet, which is reasonably well satisfied Aaltonen et al.
+We assumed that the $\Sigma_{Q}$ and $\Sigma_{Q}^{*}$ form well-separated resonances $\Gamma_{\Sigma_{Q}},\Gamma_{\Sigma_{Q}^{*}}\ll m_{\Sigma_{Q}^{*}}-m_{\Sigma_{Q}}$ relative to the mass splitting in the heavy-quark spin symmetry doublet, which is reasonably well satisfied Aaltonen et al.
 (2012); Aaij et al.
@@
 We note that the continuum $Q\to\Lambda_{Q}\pi$ background to Eq.
-(VII.2) is in fact described by a heavy-light DiFF that is of interest of its own, with the pion polar angle tracking its residual momentum fraction $z_{\pi}$ in this case.
-Importantly, however, no feature appears around $m_{\Lambda_{Q}\pi}\approx m_{\Sigma_{Q}^{*}}$ in the double-differential $(z_{\pi},m_{\Lambda_{Q}\pi})$ distribution in the background case.
+(VII.2) initiated by a boosted heavy quark is in fact described by a heavy-light DiFF that is of interest of its own, and upon expansion in $\Lambda_{\mathrm{QCD}}/m_{Q}\ll 1$ takes the same form as Eq.
+(S14) with a more differential nonperturbative matrix element $\chi_{\Lambda_{Q}\pi}$,
+$\frac{\mathrm{d}\sigma_{\Lambda_{Q}\pi X}^{\mathcal{O}_{\ell}}}{\mathrm{d}^{3}\bm{P}_{\Lambda_{Q}\pi}^{\mathrm{lab}}\,\mathrm{d}m_{\Lambda_{Q}\pi}\,\mathrm{d}\cos\theta}=\frac{\mathrm{d}\sigma_{QX}}{\mathrm{d}^{3}\bm{P}_{Q}^{\mathrm{lab}}}\,\int\!\mathrm{d}E_{\pi}\,\mathrm{d}p_{\pi}^{z}\,\chi_{\Lambda_{Q}\pi}(E_{\pi},p_{\pi}^{z})\,\delta\bigl{(}m_{\Lambda_{Q}\pi}-m_{\Lambda_{Q}}-m_{\pi}-E_{\pi}\bigr{)}\,\delta\Bigl{(}\cos\theta-\frac{p_{\pi}^{z}}{\lvert\bm{p}_{\pi}\rvert}\Bigr{)}\,,$ (S19)
+where $m_{\Lambda_{Q}\pi}$ is the invariant mass of the candidate pair and $E_{\pi}$ and $\bm{p}_{\pi}$ are the energy and three-momentum of the pion in the candidate center-of-mass frame, $\bm{p}_{\pi}=-\bm{p}_{\Lambda_{Q}}$.
+(For clarity, we have made explicit that the approximately identical kinematics $\bm{P}_{\Lambda_{Q}\pi}^{\mathrm{lab}}=\bm{P}_{Q}^{\mathrm{lab}}+\mathcal{O}(\Lambda_{\mathrm{QCD}})$ of the candidate pair and the heavy-quark progenitor are evaluated in the lab frame instead; in practice this dependence will always be integrated over a certain bin at the level of the prefactor.) The dependence of $\chi_{\Lambda_{Q}\pi}$ on $p_{\pi}^{z}$, which determines the $\cos\theta$ shape of the background, and the correlation between $p_{\pi}^{z}$ and $E_{\pi}$, satisfy the obvious on-shell constraints $E_{\pi}\geq m_{\pi}$ and $\lvert p_{\pi}^{z}\rvert\leq\sqrt{E_{\pi}^{2}-m_{\pi}^{2}}$, but are complicated in general.
+Importantly, however, nothing privileges the point $E_{\pi}\approx m_{\Sigma_{Q}^{*}}-m_{\Lambda Q}-m_{\pi}$ in the double-differential continuum fragmentation function $\chi_{\Lambda_{Q}\pi}$, and thus no feature appears around $m_{\Lambda_{Q}\pi}\approx m_{\Sigma_{Q}^{*}}$ in the double-differential $(m_{\Lambda_{Q}\pi},\cos\theta)$ distribution in the background case.
 We therefore expect that a standard sideband subtraction will remain viable to remove the background also bin by bin in $\theta$.